# Evil twins are not that evil: Qualitative insights into machine-generated prompts

### v3

@@ 1 Introduction
From a theoretical point of view, our study offers new insights into LM language processing in general. From a practical point of view, it highlights which aspects of LMs we should pay attention to, if we want to make them more robust to harmful autoprompts (or, conversely, to develop more efficient benign autoprompt generation techniques).
@@ 3 Experimental setup
We use decoder-only LMs from the Pythia (Biderman et al., 2023) and OLMo (Groeneveld et al., 2024) families, as these are fully open-source models whose training data are publicly available. Specifically, in the text we discuss the results we obtained with Pythia-6.9B, and we replicate the main experiments with Pythia-1.4B, Pythia-12B, OLMo-1B, OLMo-7B, and OLMo-7B-Instruct in Appendix A, reporting similar results.
Data collection
Prompt optimization
Data-set statistics
The final data-set we use for the Pythia-6.9B experiments reported in the main text consists of 208 triples of original prompt, autoprompt and continuation. [^5] We study relatively few autoprompts as it is very time-consuming to extract them for large model. Replicating the experiment with larger autoprompt sets using smaller models led to comparable results (Appendix A). The average original prompt length is of 39.3 tokens (s.d. 13.4); that of the continuations is of 8.4 tokens (s.d. 2.4). [^6] Datasets and code are uploadedas supplementary materials, and will be made available upon publication.
4 Experiments
4.1 Pruning autoprompts
Methodology
@@ 4.1 Pruning autoprompts
Table 1: Randomly selected examples of autoprompts and original prompts, with prunable tokens in bold. For original prompts, only the last 10 tokens are shown. ’?’ = difficult-to-render characters.
@@
Table 2: Top-10 kept or pruned tokens ranked by local mutual information for autoprompt and original prompts (for each cell, top-left has the highest value and bottom-right the lowest). Tokens are printed as follows: content word, artifact and punctuation, function word
@@
Figure 2: Counts of tokens that were pruned (dark orange) and kept (yellow) by position, where 0 is the last position. Tokens at last position are extremely unlikely to be pruned.
@@
Table 3: Number of equivalent substitutes admitted by each autoprompt token. Results are shown for all, lang-like tokens, and non-ling tokens. Legend: Avg=average; Med=median; >50:of tokens with more than 50 equivalents.
Figure 3: Pile-based log frequency distributions of bigrams in the original prompts, autoprompts and at the autoprompt/continuation boundary (ap/cont boundary). Log(0) conventionally set to -1. Red line = median; boxes span interquartile ranges.
4.2 Replacing autoprompt tokens
(a) Autoprompts
(b) Original prompts (last 10 tokens)
Figure 4: Average proportions of replacement effect types by position on pruned prompts, aligned from right (whiskers show standard deviations). Null-effect replacements leave the continuation unchanged. Moderate replacements have BLEU $\geq 0.2$. Strong replacements have BLEU $<0.2$.
(a) Autoprompts
(b) Original Prompts
@@ 4.2 Replacing autoprompt tokens
Working from now on with the pruned autoprompts, we replace the token in each position in turn with one of the 10k most frequent tokens from the Pile. We quantify the impact of the ablations in terms of BLEU score with respect to the original continuation. The ablation results are summarized in Fig. 4(a), where replacements are binned based on the impact they have on the continuation (examples are presented in the tables of Appendix B).
@@
First, we confirm that non-pruned tokens in all positions play a significant role in generating the continuation, as shown by the fact that most replacements have a strong impact on BLEU. However, for all positions except the last, we also see that a non-negligible proportion of replacements do not affect the continuation at all, and in a significant proportion of cases the continuation is only mildly affected (as the examples in Table 11 of Appendix B show, even a BLEU score around 0.2 typically corresponds to a continuation that is quite similar to the original). We confirm moreover the special role of the last token, that can almost never be replaced without catastrophic results. More generally, as we approach the last position, it is increasingly more difficult to find replacements that do not strongly affect the continuation.
@@
Table 4: Examples of replacement for autoprompts and original prompts (10 last tokens). Color represents the number of substitutes: $>50$ substitutes and $<50$ substitutes. When there are less than 50 substitutes, a random subset is displayed. Difficult-to-render characters are replaced by ’?’. More examples are provided in Table B (appendix).
@@
We further looked at the equivalent sets, consisting, for a given token, of the substitutes that keep the continuation unchanged. Table 3 shows $76.4\%$ of the tokens can be substituted by at least one equivalent, and each token has $302.8$ substitutes on average. But the size of the equivalent set is highly variable, with half of the tokens having $15.0$ or fewer equivalents. This disparity is intuitively informative about the role of each token. In particular, we identify the presence of fillers that can be replaced by a large number of substitutes, and key tokens, that admit more restrained equivalent sets and must thus carry more specific information. This is evidenced by the fact that language-like tokens tend to admit less equivalent substitutes than non-linguistic tokens. Table 4.2 (top section) shows examples of autoprompts with tokens color-coded based on their equivalent set size. When the equivalent set is small, we provide a random sample of it as an illustration.
@@
We measured the semantic consistency within equivalent sets in order to determine whether substitution in autoprompt is governed by semantic similarity (akin to synonymy in natural language) or by other factors. Specifically, we used FastText (Bojanowski et al., 2017) to compute the average semantic similarity: (a) between a token and each of its equivalents; and (b) among the equivalents themselves (including the original token). [^9] For technical reasons, due to the difficulty of obtaining reliable representations for non-linguistic tokens, we focused only on language-like tokens. The first measure indicates how semantically relevant the substitutes are in relation to the original token, while the second measure reflects the size of the substitute semantic space. A value close to $1$ suggests that the substitutes cluster within a small semantic region, meaning they are semantically consistent with one another. We plot semantic similarity against set size in Fig 5(a). As expected, small equivalent sets –corresponding to key tokens– tend to be semantically coherent on average. In this sense, they approximate near synonyms in natural language. This is confirmed by the examples in Table 4.2, e.g., ‘Ireland’ is replaced by ‘Irish’ and ‘Tehran’ by ‘Iran’. Nonetheless, the semantic relation is often approximate, e.g., ‘before’ being replaced by ‘then’, ‘change’ by ‘subsequently’ or ‘forming’ by ‘meeting’, ‘leaving’, ‘remaining’, etc. In contrast, large sets –associated to fillers– often include tokens that are semantically unrelated to each other. We show in Table 5 a subset of the most frequently found tokens in large sets. It mostly consists of subwords (sometimes interpretable morphemes), digits and fragments of named entities that do not carry strong semantic information.
105 <EOT> () 113 ini Mar ran 25 be Green 101 56 ¸ mat ato ator Rock 35 115 Moore 114 SS lim eter ester abe uns Head Char ? ? ? ac aa aur ? ong ost den FF 54 a ab ud sl rim st ? 022 1 37 91 INT op ore ef rel cha ast ives
Table 5: List of the $60$ tokens most frequently found in large equivalent sets (observed in about $10\%$ of the sets). Difficult-to-render characters are replaced by ’?’. Tokens are marked as follows: digit, artifact and punctuation, named entity, subword and morphological unit. See more in Table 14.
@@
The four authors of the paper manually annotated the autoprompt dataset with a binary label marking, for each token, if it is intuitively important or not for the generation of the continuation. The annotation was made a posteriori, using the full autoprompt and the generated continuation (see appendix C for a description of the annotation process). Results show that autoprompts possess interpretable properties, as the labeling is correlated with the number of replacements a token might have. Indeed, we measured that the median size of the equivalent set for tokens deemed important is $2$, against $15$ for those deemed replaceable, with the effect observed both for language-like and non-linguistic tokens (cf. Table 16).
@@
In cases where the replacement causes only a moderate change in the continuation, we see signs of “compositionality,” in the sense that the continuation only displays a few new tokens broadly reflecting the meaning of the replacement. Some examples are presented in Table 6. We make the intuition more quantitative as follows. First, to facilitate automated similarity analysis, we extracted all cases where the replacement leads to the change of a single (typographic) word in the continuation (about 3% of the total). For these cases, we used FastText to measure the semantic similarity of both the original autoprompt token and its replacement to the original word in the continuation and to the changed one. We found that the original token is more similar to the new continuation word (vs. the original one) in only 48% of the cases, whereas the replacement token is more similar to the new continuation in 59% of the cases. We thus conclude that, indeed, there is a tendency for at least this type of replacement to work compositionally, with a small change in the autoprompt leading to a semantically consistent change in the continuation. This, in turn, suggests that autoprompts do not function as unanalyzable holistic wholes. Their “meaning” to the model derives, at least partially, from assembling the meaning of its parts, as with natural language sequences. However, this looks nothing like the one performed by natural language syntax.
@@
Table 6: Example autoprompt token replacements leading to a small, interpretable change in the continuation (legend: replaced/replacement in autoprompt; and original/new in continuation).
@@ 4.3 Shuffling autoprompt tokens
The average BLEU when shuffling all tokens is at 0.03 (s.d. 0.04) and at 0.06 (s.d. 0.11) when leaving the last token in its slot. This difference is highly significant (paired t-test, $p<0.001$). [^10] The difference stays comparably significant if, in the first condition, we leave a random non-last token fixed, so that the same number of tokens is shuffled in the two cases. However, the low BLEU values suggest that, contrary to our conjecture, the autoprompt tokens before the last are not a bag of keywords, since their order matters as well. One possibility is that, while autoprompts as a whole do not constitute syntactically well-formed sequences, they are composed of tight sub-sequences that should not be separated. For example, given that modern tokenizers split text at the sub-word level, token-level shuffling will arbitrarily break words. Some support for the view that the catastrophic effect of shuffling pre-last tokens is due to short-distance dependencies comes by looking at the cases in which a bigram in an autoprompt (excluding the last position) is also attested in the Pile corpus, either in the original or in the inverted order. In 60.5% of these cases, the Pile frequency of the original bigram is larger than that of the inverted one, suggests some degree of natural local ordering among autoprompt tokens.
@@ 4.4 Making human prompts more autoprompt-like
We replicate the token-replacement experiment on the pruned original prompts, obtaining the results summarized in Figure 4(b). Again, tokens become more replaceable as we move away from the end of the prompt, confirming the crucial role played by the very last token. Moreover, we confirm the presence of both key tokens, having few semantically related substitutes, and fillers with numerous semantically inconsistent substitutes (Figure 5(b)). Table 4.2 (bottom) shows examples in which the original prompt, despite pruning and replacement among the last 10 tokens, still triggers the same continuation. We see how the same principles that might explain the success of autoprompts are at work here, suggesting how autoprompts might take shape during the induction process.
@@ 5 Discussion
Our findings about autoprompts, confirmed by autoprompt-inspired ablations of natural prompts, suggest that LMs might rely on a simplified model of language, where not all tokens have specific syntactic and semantic functions in an abstract syntactic tree. We note that the phenomenon of relying on over-simplified representations of the data is not specific to LMs. Convolutional Neural Network classifiers of visual data also latch onto superficial correlations in the data, leading to poor ood generalization (Jo & Bengio, 2017; Ilyas et al., 2019; Yin et al., 2019; Geirhos et al., 2020).
@@
Ethics Statement
If we do not achieve a genuine understanding of how LMs process and generate text, we cannot fully control their behavior and mitigate unintended or intentional harm. Opaque autoprompts are an indication that there are important aspects of LM prompting and generation that are still out of our control. Our investigation into the nature of this phenomenon contributes to a better understanding of how LMs work and, thus, ultimately, to make them safer and more predictable.
@@ Appendix A Results with other models
The results of the autoprompt pruning and shuffling experiments are presented in Table 8. For all models, there is a difference in BLEU when shuffling all tokens vs. keeping last token fixed (paired t-test significant at p < 0.01). Token pruning distribution by position is shown in Figures 17 18 19 20 21 (a).
@@
For OLMo models, we estimate the top 10k most frequent tokens to be used in the replacement experiments using a sample of approximately 10 billion tokens from the Dolma corpus, which was used to train this model. (Soldaini et al., 2024). Proportions of replacement effect type by position are reported in Figures 17 18 19 20 21 (b).
@@
Similarly, we measure the semantic consistency of the equivalent sets in Figures 17 18 19 20 21 (c).
@@ A.1 Making prompts more autoprompt-like
The results of the pruning and shuffling experiments of natural prompts are presented in Table 9. For all models except OLMo-7B-Instruct, there is a difference in BLEU when shuffling all tokens vs. keeping last token fixed (paired t-test significant at p < 0.01). Token pruning by position is reported in Figures 17 18 19 20 21 (a).
@@
Proportions of replacement effect type by position are reported in Figures 17 18 19 20 21 (b).
@@
Semantic consistency of the equivalent sets are reported in Figures 17 18 19 20 21 (c).

### v4

@@ Abstract
Evil twins are not that evil:
Qualitative insights into machine-generated prompts
Nathanaël Carraz Rakotonirina* 1, Corentin Kervadec* 1, Francesca Franzon1, Marco Baroni1,2 1Universitat Pompeu Fabra, Barcelona 2ICREA, Barcelona
@@ 1 Introduction
From a theoretical point of view, our study offers new insights into LM language processing in general. From a practical point of view, it highlights which aspects of LMs we should pay attention to, if we want to make them more robust to harmful autoprompts (or, conversely, to develop more efficient benign autoprompt generation techniques). We present the first thorough analysis of opaque machine-generated prompts, or autoprompts, pertaining to 6 LMs of different sizes and families, focusing on minimal pairs of natural language prompts and their “evil twins”, i.e., opaque autoprompts that lead to the same continuation.
@@ 3 Experimental setup
We use decoder-only LMs from the Pythia (Biderman et al., 2023) and OLMo (Groeneveld et al., 2024) families, as these are fully open-source models whose training data are publicly available. Specifically, in the text we discuss the results we obtained with Pythia-6.9B, and we replicate the main experiments with Pythia-1.4B, Pythia-12B, OLMo-1B, OLMo-7B, and OLMo-7B-Instruct in App. A, reporting similar results.
gray!50
@@
Table 1: Randomly selected examples of autoprompts and original prompts for Pythia-6.9B, with prunable tokens in bold. For original prompts, only the last 10 tokens are shown. ’?’ = difficult-to-render characters.
black
gray!50
Table 2: Top-10 kept or pruned tokens for Pythia-6.9B, ranked by local mutual information for autoprompt and original prompts (for each cell, top-left has the highest value and bottom-right the lowest). Tokens are printed as follows: content word, artifact and punctuation, function word
black
Data collection
Prompt optimization
Data-set statistics
The final data-set we use for the Pythia-6.9B experiments reported in the main text consists of 208 triples of original prompt, autoprompt and continuation. [^5] We study relatively few autoprompts as it is very time-consuming to extract them for large model. Replicating the experiment with larger autoprompt sets using smaller models led to comparable results (App. A). The average original prompt length is of 39.3 tokens (s.d. 13.4); that of the continuations is of 8.4 tokens (s.d. 2.4). [^6] Datasets and code are uploadedas supplementary materials, and will be made available upon publication.
4 Experiments
4.1 Pruning autoprompts
Methodology
@@ 4.1 Pruning autoprompts
Figure 2: Counts of tokens that were pruned (dark orange) and kept (yellow) by position for Pythia-6.9B, where 0 is the last position. Tokens at last position are extremely unlikely to be pruned.
@@
Figure 3: Pile-based log frequency distributions of bigrams in the original prompts, autoprompts and at the autoprompt/continuation boundary (ap/cont boundary) for Pythia-6.9B. Log(0) conventionally set to -1. Red line = median; boxes span interquartile ranges.
4.2 Replacing autoprompt tokens
(a) Autoprompts
(b) Original prompts (last 10 tokens)
Figure 4: Average proportions of replacement effect types by position on pruned prompts for Pythia-6.9B, aligned from right (whiskers show standard deviations). Null-effect replacements leave the continuation unchanged. Moderate replacements have BLEU $\geq 0.2$. Strong replacements have BLEU <0.2<0.2.
@@ 4.2 Replacing autoprompt tokens
Table 3: Number of equivalent substitutes admitted by each autoprompt token for Pythia-6.9B. Results are shown for all, lang-like tokens, and non-ling tokens. Legend: Avg=average; Med=median; >50:% of tokens with more than 50 equivalents.
@@
Working from now on with the pruned autoprompts, we replace the token in each position in turn with one of the 10k most frequent tokens from the Pile. We quantify the impact of the ablations in terms of BLEU score with respect to the original continuation. The ablation results are summarized in Fig. 4(a), where replacements are binned based on the impact they have on the continuation (examples are presented in App. B).
@@
First, we confirm that non-pruned tokens in all positions play a significant role in generating the continuation, as shown by the fact that most replacements have a strong impact on BLEU. However, for all positions except the last, we also see that a non-negligible proportion of replacements do not affect the continuation at all, and in a significant proportion of cases the continuation is only mildly affected (as the examples in Table 10 of App. B show, even a BLEU score $\approx$ 0.2 typically corresponds to a continuation that is quite similar to the original). We confirm moreover the special role of the last token, that can almost never be replaced without catastrophic results. In general, as we approach the last position, it is increasingly more difficult to find replacements that do not strongly affect the continuation.
gray!50
@@
Table 4: Examples of replacement for autoprompts and original prompts (10 last tokens) for Pythia-6.9B. Color represents the number of substitutes: >50>50 substitutes and <50<50 substitutes. When there are less than 50 substitutes, a random subset is displayed. Difficult-to-render characters are replaced by ’?’. More examples in Table 12 (appendix).
black
@@
We further looked at the equivalent sets, consisting, for a given token, of the substitutes that keep the continuation unchanged. We measure that $76.4\%$ of the tokens can be substituted by at least one equivalent, and each token has $302.8$ substitutes on average (cf. Table  3). But the size of the equivalent set is highly variable, with half of the tokens having $15.0$ or fewer equivalents. This disparity is intuitively informative about the role of each token. In particular, we identify the presence of fillers that can be replaced by a large number of substitutes, and key tokens, that admit more restrained equivalent sets and must thus carry more specific information. This is evidenced by the fact that language-like tokens tend to admit less equivalent substitutes than non-linguistic tokens. Table 4 (top section) shows examples of autoprompts with tokens color-coded based on their equivalent set size. When the equivalent set is small, we provide a random sample of it as an illustration.
@@
We measured the semantic consistency within equivalent sets in order to determine whether substitution in autoprompt is governed by semantic similarity (akin to synonymy in natural language) or by other factors. Specifically, we used FastText (Bojanowski et al., 2017) to compute the average semantic similarity: (a) between a token and each of its equivalents; and (b) among the equivalents themselves (including the original token). [^9] For technical reasons, due to the difficulty of obtaining reliable representations for non-linguistic tokens, we focused only on language-like tokens. The first measure indicates how semantically relevant the substitutes are in relation to the original token, while the second measure reflects the size of the substitute semantic space. A value close to $1$ suggests that the substitutes cluster within a small semantic region, meaning they are semantically consistent with one another. We plot semantic similarity against set size in Fig 5(a) (App. B). As expected, small equivalent sets –corresponding to key tokens– tend to be semantically coherent on average. In this sense, they approximate near synonyms in natural language. This is confirmed by the examples in Table 4, e.g., ‘Ireland’ is replaced by ‘Irish’ and ‘Tehran’ by ‘Iran’. Nonetheless, the semantic relation is often approximate, e.g., ‘before’ being replaced by ‘then’, ‘change’ by ‘subsequently’ or ‘forming’ by ‘meeting’, ‘leaving’, ‘remaining’, etc. In contrast, large sets –associated to fillers– often include tokens that are semantically unrelated to each other. We show in Table 13 (App. B) the most frequently found tokens in large sets, mostly consisting of subwords (sometimes interpretable morphemes), digits and fragments of named entities that don’t carry strong semantic information.
@@
The four authors of the paper manually annotated the autoprompt dataset with a binary label marking, for each token, if it is intuitively important or not for the generation of the continuation. The annotation was made a posteriori, using the full autoprompt and the generated continuation (see app. C for a description of the annotation process). Results show that autoprompts possess interpretable properties, as the labeling is correlated with the number of replacements a token might have. Indeed, we measured that the median size of the equivalent set for tokens deemed important is $2$, against $15$ for those deemed replaceable, with the effect observed both for language-like and non-linguistic tokens (Table 15).
@@
In cases where the replacement causes only a moderate change in the continuation, we see signs of “compositionality,” in the sense that the continuation only displays a few new tokens broadly reflecting the meaning of the replacement. Some examples are presented in Table 5. We make the intuition more quantitative as follows. First, to facilitate automated similarity analysis, we extracted all cases where the replacement leads to the change of a single (typographic) word in the continuation (about 3% of the total). For these cases, we used FastText to measure the semantic similarity of both the original autoprompt token and its replacement to the original word in the continuation and to the changed one. We found that the original token is more similar to the new continuation word (vs. the original one) in only 48% of the cases, whereas the replacement token is more similar to the new continuation in 59% of the cases. We thus conclude that, indeed, there is a tendency for at least this type of replacement to work compositionally (a small change in the autoprompt leads to a semantically consistent change in the continuation). This, in turn, suggests that autoprompts do not function as unanalyzable holistic wholes. Their “meaning” to the model derives, at least partially, from assembling the meaning of its parts, as with natural language sequences. However, this looks nothing like the one performed by natural language syntax.
gray!50
@@
Table 5: Example autoprompt token replacements leading to a small, interpretable change in the continuation for Pythia-6.9B (legend: replaced/replacement in autoprompt; and original/new in continuation).
black
@@ 4.3 Shuffling autoprompt tokens
The average BLEU when shuffling all tokens is at 0.03 (s.d. 0.04) and at 0.06 (s.d. 0.11) when leaving the last token in its slot. This difference is highly significant (paired t-test, $p<0.001$). However, the low BLEU values suggest that, contrary to our conjecture, the autoprompt tokens before the last are not a bag of keywords, since their order matters as well. One possibility is that, while autoprompts as a whole do not constitute syntactically well-formed sequences, they are composed of tight sub-sequences that should not be separated. For example, given that modern tokenizers split text at the sub-word level, token-level shuffling will arbitrarily break words. Some support for the view that the catastrophic effect of shuffling pre-last tokens is due to short-distance dependencies comes by looking at the cases in which a bigram in an autoprompt (excluding the last position) is also attested in the Pile corpus, either in the original or in the inverted order. In 60.5% of these cases, the Pile frequency of the original bigram is larger than that of the inverted one, suggests some degree of natural local ordering among autoprompt tokens.
@@ 4.4 Making human prompts more autoprompt-like
We replicate the token-replacement experiment on the pruned original prompts, obtaining the results summarized in Figure 4(b). Again, tokens become more replaceable as we move away from the end of the prompt, confirming the crucial role played by the very last token. Moreover, we confirm the presence of both key tokens, having few semantically related substitutes, and fillers with numerous semantically inconsistent substitutes (Figure 5(b)). Table 4 (bottom) shows examples in which the original prompt, despite pruning and replacement among the last 10 tokens, still triggers the same continuation. We see how the same principles that might explain the success of autoprompts are at work here, suggesting how autoprompts might take shape during the induction process.
@@ 5 Discussion
Our findings about autoprompts, confirmed by autoprompt-inspired ablations of natural prompts, suggest that LMs might rely on a simplified model of language, where not all tokens have specific syntactic and semantic functions in an abstract syntactic tree. We note that the phenomenon of relying on over-simplified representations of the data is not specific to LMs. Convolutional Neural Network classifiers of visual data also latch onto superficial correlations in the data, leading to poor ood generalization (Jo and Bengio, 2017; Ilyas et al., 2019; Yin et al., 2019; Geirhos et al., 2020).
@@ Appendix A Results with other models
The results of the autoprompt pruning and shuffling experiments are presented in Table 7. For all models, there is a difference in BLEU when shuffling all tokens vs. keeping last token fixed (paired t-test significant at p < 0.01). The difference stays comparably significant if, in the first condition, we leave a random non-last token fixed, so that the same number of tokens is shuffled in the two cases. Token pruning distribution by position is shown in Figures 6 7 8 9 10 (a).
@@
For OLMo models, we estimate the top 10k most frequent tokens to be used in the replacement experiments using a sample of approximately 10 billion tokens from the Dolma corpus, which was used to train this model. (Soldaini et al., 2024). Proportions of replacement effect type by position are reported in Figures 6 7 8 9 10 (b).
@@
Similarly, we measure the semantic consistency of the equivalent sets in Figures 6 7 8 9 10 (c).
@@ A.1 Making prompts more autoprompt-like
The results of the pruning and shuffling experiments of natural prompts are presented in Table 8. For all models except OLMo-7B-Instruct, there is a difference in BLEU when shuffling all tokens vs. keeping last token fixed (paired t-test significant at p < 0.01). Token pruning by position is reported in Figures 6 7 8 9 10 (a).
@@
Proportions of replacement effect type by position are reported in Figures 6 7 8 9 10 (b).
@@
Semantic consistency of the equivalent sets are reported in Figures 6 7 8 9 10 (c).